OpenStudy (anonymous):

Rewrite the function in the form y = ( x − h ) 2 + k, and find the vertex: y = x2 − 12x A. y = (x + 6 ) 2 − 36, vertex = ( 2, −6 ) B. y = ( x − 6 ) 2 + 36, vertex = ( 6, 0 ) C. y = ( x − 6 ) 2 − 36, vertex = ( 6, −36 ) D. y = x 2 − 36, vertex = ( −6, −6 )

Parth (parthkohli):

Do you know how to complete the square?

Parth (parthkohli):

@jordanguruu ?

OpenStudy (anonymous):

no

Parth (parthkohli):

satellite = jordanguruu?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

i only know how to compute \(-\frac{b}{2a}=-\frac{-12}{2}=6\) so vertex is \((6,-36)\) and it looks like \(y=(x+6)^2-36\)

Parth (parthkohli):

In any quadratic equation, you must add (\({b \over 2})^2\) to both sides.

OpenStudy (anonymous):

i still do not understand

OpenStudy (anonymous):

can you help me?

OpenStudy (anonymous):

yes... but the answer is already given.... do you need furthur explanation?

OpenStudy (anonymous):

i dont see the answer

OpenStudy (anonymous):

lemme check...

OpenStudy (anonymous):

yes... you are correct...

OpenStudy (anonymous):

ok.. the x coordinate of the vertex of any parabola \[\large y=ax^2+bx+c\] is obtained by -b/(2a) where a , b, and c are the coefficients of the quadratic equation.

OpenStudy (anonymous):

still with me?

OpenStudy (anonymous):

i think so so the answer would be C? A. y = (x + 6 ) 2 − 36, vertex = ( 2, −6 ) C. y = ( x − 6 ) 2 − 36, vertex = ( 6, −36 )

OpenStudy (anonymous):

correct... C... :)

Parth (parthkohli):

Yes!

Parth (parthkohli):

A. doesn't have the vertex written correctly

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

good work man... :)

OpenStudy (anonymous):

i have one more question can you help me?

OpenStudy (anonymous):

post it up as a new question...

OpenStudy (anonymous):

Use the quadratic formula to solve the equation: 6x 2 − 5x + 1 = 0 A. x = 1 or x = – 1⁄2 B. x = 2 or x = −1 C. x = – 1 ⁄ 3 or x = 3 D. x = 1 ⁄ 2 or x = 1 ⁄ 3

OpenStudy (anonymous):

can you help?